Diffusion Morning Report 06/29/11

Ted Berg tberg at stanford.edu
Wed Jun 29 06:21:42 PDT 2011


Good morning All,
     Looking good this morning. Epi had HCL flow issues. TEOS tube 
changed- going back together. LTO had low dep rate. Thermco 1 low growth 
rate-adjusting door. Thermco 2 no tlc flow. Thermco Nitride higher 
pressure during dep.


					
	Date: 	6/29/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	up 				
Tylan 2 	up 				
Tylan 3 	up 				
Tylan 4 	up 				
Tylan 5-BBr3 	up 				
Tylan 6-POCl3 	up 				
Tylan 7-TEOS2 	yellow 	tube change 	0/30 		2
Tylan 8-SiGe 	up 		1.94/30 		
Tylan 9-Poly 	up 		22.3/30 		
Tylan 10-Nitride 	up 		16.51/30 		
Tylan 11-LTO 	yellow 	low dep rate 	10.4/30 		
Tylan 12-FGA 	up 				
Thermco 1 	yellow 	low growth rate- Adjusted Door 			
Thermco 2 	yellow 	no tlc flow 			
Thermco 3 	future 				
Thermco4 	future 				
Thermco Poly1 	up 	
			
	up 				
Thermco Nitride1 	yellow 	pressure issues 			
Thermco bpsg 	up 				
AMAT Epi 	yellow 	HCL flow 			1
TEL Spa 	up 				
Burn Box 	up 				
Ebeam 	down 	cryo compresssor dead 			
ZYGO 	up 	
			
FGA2 	up 				
SiH4 for furnaces 		B on line pressure 938 psi 	changeover working well		


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110629/0d5ac204/attachment.html>


More information about the diffusion mailing list